OpenStudy (anonymous):

Rotate the region bounded by the given curves about the given line and find the volume resulting solid. x=1+y^2 and the xaxis y=1 to y=2 about the xaxis

OpenStudy (anonymous):

 \[f(y)=\int\limits\limits_{1}^{2}(1+y^{2})dy=y+\frac{y^3}{3}/_{1}^{2} =(2+\frac{8}{3})-(1+\frac{1}{3})=3.333Unit^{2}\] Circumference of circle that the rotating shape diameter = 2 \[Circumference=2\times \pi = 6.28 unit\] volume = 6.28 Unit × 3.333 unit^2 = 20.93124 unit ^3 you are welcome .